# Ultra-Fast Language Generation via Discrete Diffusion Divergence Instruct

### v1

@@ F.6 Protein Sequence Generation
Figure 5: pLDDT comparison between DiDi-Instruct and the DPLM-150M across different sequence lengths ($L=100,200,300,400,500$) and numbers of function evaluations (NFEs). Our method consistently outperforms the teacher, achieving up to +10 pLDDT gains at shorter sequence lengths (e.g., $L=100$) and maintaining superior structural confidence across all lengths, even with substantially fewer sampling steps. Moreover, the distilled student exhibits more stable performance across NFEs, whereas the teacher shows larger variability as the number of steps increases.
Dataset and Evaluation Metric.
The UniRef50 dataset (Suzek et al., 2014) contains approximately 45 million protein sequences, comprising roughly 14 billion amino acid tokens. Following previous work (Wang et al., 2024), we adopt a vocabulary of 33 tokens and a maximum sequence length of 1024, with longer proteins chunked into shorter chunks. We compute the predicted local distance difference test (pLDDT) score using the ESMFold model (Lin et al., 2023) to assess the foldability of the generated protein sequences, with values above 70 indicating high structural confidence. We report pLDDT score across different sequence lengths and varying numbers of function evaluations (NFEs).
Experiment Details.
Our model architecture follows the DPLM-150M setting (Wang et al., 2024), with the student model and discriminator matched in size, while the discriminator incorporates a modified prediction head that outputs a scalar value. We set the learning rate of the student model to $1\times 10^{-5}$ and that of the discriminator to $5\times 10^{-6}$, with 1000 warm-up steps for each. All other training hyperparameters and sampling strategies follow those used in the text modeling task. The distillation of DPLM was completed in approximately two hours on a single H100 GPU, using batches of up to 4096 tokens.
Protein Sequence Examples.
We provide visualizations of generated protein sequences to compare structural plausibility between the teacher and our distilled model. Figure 6 shows low-confidence outputs from the teacher under limited NFEs, while Figure 7 presents high-confidence examples produced by DiDi-Instruct in Figure 7.
@@
Figure 7: Visualization examples of high-confidence protein sequences ($\text{pLDDT}>70$) generated by the DiDi-Instruct model, with lengths ranging from 50 to 500 across different numbers of function evaluations (NFEs).
@@ F.7 Text examples
Fluency and Repetition.
All models (including the teacher and students) generally produce fluent and grammatically correct sentences. The most significant exception is the 8-step student, which suffers from severe phrasal repetition (e.g., repeating “affirmative action program"), a common artifact of extreme model compression. This issue is effectively eliminated in student models with 16 NFEs or more, which exhibit fluency comparable to the teacher.
Coherence and Topic Adherence.
This is where the most critical differences emerge. The 1024-step teacher struggles with global coherence, frequently shifting between unrelated topics within a single output. In contrast, while the 8-step student is incoherent due to repetition, the 16-step student already demonstrates stronger paragraph-level topic adherence than the teacher. This ability to maintain a consistent narrative thread strengthens progressively. The 128-step student excels at this, developing a central theme with supporting details over a longer text, showcasing high global coherence.
Informativeness and Specificity.
The informativeness of the generated text correlates strongly with coherence. The teacher’s outputs, despite containing specific details, are uninformative as a whole due to topic shifts. The 8-step student’s text has very low information content due to repetition. As NFEs increase from 16 to 128, the students’ generations become increasingly specific and informative. For example, the 64-step model constructs a detailed news-style report, and the 128-step model builds a multi-faceted story, both rich in contextual details.
@@
Text obtained with $1024$ NFEs teacher model (1/3). Perplexity=$40.48$, Entropy=$5.22$. He was mid-morning early, and I started forming a motorcade for him. He stopped every NYPD car, and I radioed back to every bicycle officer that was in there, to protect him. I did not hear you said to be involved in it. That you were involved with that campaign. I never asked anybody to test it. You did need help you leaked it, not made it for yourself. He parked and it was then that I saw that president and ended up there, speaking with a journalist. The Prince was granted anonymity and noted that the FBI was aware of the familiarity with Russia and was conducting the investigation as appropriate. In some accounts, Casey Wilson, senior counsel for the Bonden Bisson, told The Post that the President is aware of the inquiry. But Mr. Bakers testimony is the first time someone has long taken something of concern into the hands of their officials probing Trumps personal relationships with foreign officials. Former FBI regional director James Kogan, a former officer who met with President Trump, was once briefed about the inquiry from the Russian District Attorney General. Victor Atkinson, an FBI chief-officer who is retired, hacked the personal PINs and cell phone lines of the two Trump officials to conceal the information, as shown inside the countrys electronic networks. Anonymous messages can appear in our apps like Twitter and Facebook and if ones political actions resists they become blocked. The Daily Signal columnist, and former Edward Snowden contractor, last summer. According to the piece he leaked, his company was set up to log phone calls for the officials so that some of Hillarys callers would end up because Trump and businesses required a gag. I was using this same setup last week during my private conversation with Trump. I called him after this piece of political reporting and "faming," he said of Barack Obama, as he told me last May. The conversation about the leak was livestreamed at 1 p.m. this evening. CAIRCLE MEDALE, RAC.Y. Authorities a 28-year-old Burnet County receiving a call at 10:30 p.m. live across the street in Greene County reported that the man he heard was telling a neighbor to kill him. Abedoy and Terry Angelo, 24, found the house. They were charged with disturbing the peace, said Greene Police Detective Nathaniel Clay. "There is a door with a phone in a glass covering, and there is a grenade at the window," Clay said. "The caller smashed out the window, indicating it was the mans father." Hes had a similar house for a week. They were just found. Police said Angelo said he thought up the name while he was delivering. "Police determined that the most recent break in the home had been set in place," Clay said. He said assistance.
@@
Text obtained with $1024$ NFEs teacher model (2/3). Perplexity=$40.48$, Entropy=$5.22$. Montana State Supreme Court will vote Monday-2 to extend the reach of the case in Indiana until the league has their final ruling in court on November 4. The order in case reads as the first rule that the league will either settle trial court or the trial. With the lower court denying it. This means that when the ramifications stand, all players could be considered at "the game" like mayow or receive other incentives, likely a winning boost after Fridays game in Chicago. Gomez is 6-foot-6 by weight when he opens up. And probably is the future top free agent, but no players can return for taking a portion of those benefits. S. Thibodeau/AP Westwood is the standard locale to that end. Under the rules, the league must travel a minimum of four minutes per day to represent Finals 2 and 4. To every benefit from all the games, and that only players who participate play each others for each game. The league, which determines the future status of the highest free agent, will head to court in Westwood by Wednesday to avoid the headaches. The league serves as trial court in the case -- presumably with the court, which then gets to decide on any procedural cases that may arise -- before June 18, although the league says its not likely. The current clause disallows the NBA from linking any agent to the league. And prohibits it from accepting unany sponsorships that may have affected players without the approval of a director. Through its appeal, the NBA believes the Westwood will decide whether to receive any. It will update the league as it makes its decision the other day. If results are not appealed, the league ordered the court in it its ruling on March 29. It considered for an appeals court ruling on Monday, but that order has not been taken up by courts. The case seemed to already be dead Friday. After all of the courts first 18 officials handed the briefcase and could not earn a reach ruling in court itself, but it argued that it was hearing their appeal. The rule, when it is written, sends the 16-appointed back to a trial. And if they appeals the decision by following Friday to sea next Tuesday, hopefully the dispute will then be determined by the trial court. If the case is settled, the success of the Indiana incident in September could determine if the trial court can use its argument against the league and the NBA. Once both sides have signed a settlement agreement. The league says that under Indiana law the league must dismiss rulings Nov. 4 by the lower court, meaning that if it claims that a trial ruling will not be valid, it will be too late to qualify.
@@
Text obtained with $1024$ NFEs teacher model (3/3). Perplexity=$40.48$, Entropy=$5.22$. And even less because to me the LGBT community are homophobic. Were really unforgressive because generally part of us is about equality. And I guess people really do not know, if were LGBT and were not going to pay for it, wouldnt we just do it? Well, yeah, yeah, yeah, that would be great. No, I mean realized we would pay for that because then youd just have a really bad time. If something is seriously important to you and wanting to be able to participate participate in it, youd be great at that. What do you think would be of stepping up to support your local team? Is there any way your parents would have to be able to if you played to the next level. Do they feel that responsibility as parents? Yes, but I do think that when you grow up in an LGBT environment, a lot of stuff is going to happen. Thats why it has been about anything that is civil. Look, I think that its working, but obviously a family can never get through that hierarchy here because younger people are very well theyre raised, and theyre going through a lot of bad experiences around nobody that are parents. Growing up from a parent figure is just a less enjoyable environment to be able to have. We are telling the story here, so you have an answer for it. How long have you and your band have toured a college district? When were you in college? Hunsheeling Foo since I was in college. Ive never really thought about recruiting for a high school. But it wasnt just just me going onto seeing whatever Neil did or why Google just made this about it all, or it was me going into every second week. Or going into at least some draft, though, maybe thats different than knowing what this thing is. Thats why I scholarship, that made the sense. I only broke records for two years at Kansas State and Kansas, I graduated and I had many great battles afterward, so theyre always in discussion. And I have a friend recruited as a part of the Michigan teams and Michigan. David Denselmoss, former Buffalo Bills backup QB for the Eagles that coach Bruce Allen used in recruiting from Michigan for Colorado. (He was originally planned to have been a Michigan transfer.) We happened to start your press conference because you mentioned high-profile recruits but not selected in that one, either. Dolting Peps: No question. Its certainly not an uncommon one to make that mistake. But if you do ask me theres nothing I can do wrong with being among the 60 worst coaches in football.

### v2

@@ F.6 Protein Sequence Generation
Figure 5: pLDDT comparison between DiDi-Instruct and the DPLM-150M across different sequence lengths ($L=100,200,300,400,500$) and NFEs. Our method consistently outperforms the teacher, achieving up to +10 pLDDT gains at shorter sequence lengths (e.g., $L=100$) and maintaining superior structural confidence across all lengths, even with substantially fewer sampling steps. Moreover, the distilled student exhibits more stable performance across NFEs, whereas the teacher shows larger variability as the number of steps increases.
Dataset and evaluation metric. The UniRef50 dataset (Suzek et al., 2014) contains approximately 45 million protein sequences, comprising roughly 14 billion amino acid tokens. Following previous work (Wang et al., 2024), we adopt a vocabulary of 33 tokens and a maximum sequence length of 1024, with longer proteins chunked into shorter chunks. We compute the predicted local distance difference test (pLDDT) score using the ESMFold model (Lin et al., 2023) to assess the foldability of the generated protein sequences, with values above 70 indicating high structural confidence. We report pLDDT score across different sequence lengths and varying NFEs.
Experiment details. Our model architecture follows the DPLM-150M setting (Wang et al., 2024), with the student model and discriminator matched in size, while the discriminator incorporates a modified prediction head that outputs a scalar value. We set the learning rate of the student model to $1\times 10^{-5}$ and that of the discriminator to $5\times 10^{-6}$, with 1000 warm-up steps for each. All other training hyperparameters and sampling strategies follow those used in the text modeling task. The distillation of DPLM was completed in approximately two hours on a single H100 GPU, using batches of up to 4096 tokens.
Protein sequence examples. We provide visualizations of generated protein sequences to compare structural plausibility between the teacher and our distilled model. Figure 6 shows low-confidence outputs from the teacher under limited NFEs, while Figure 7 presents high-confidence examples produced by DiDi-Instruct in Figure 7.
@@
Figure 7: Visualization examples of high-confidence protein sequences ($\text{pLDDT}>70$) generated by the DiDi-Instruct model, with lengths ranging from 50 to 500 across different NFEs.
@@ F.7 Text Examples
Fluency and repetition. All models (including the teacher and students) generally produce fluent and grammatically correct sentences. The most significant exception is the 8-step student, which suffers from severe phrasal repetition (e.g., repeating “affirmative action program”), a common artifact of extreme model compression. This issue is effectively eliminated in student models with 16 NFEs or more, which exhibit fluency comparable to the teacher.
Coherence and topic adherence. This is where the most critical differences emerge. The 1024-step teacher struggles with global coherence, frequently shifting between unrelated topics within a single output. In contrast, while the 8-step student is incoherent due to repetition, the 16-step student already demonstrates stronger paragraph-level topic adherence than the teacher. This ability to maintain a consistent narrative thread strengthens progressively. The 128-step student excels at this, developing a central theme with supporting details over a longer text, showcasing high global coherence.
Informativeness and specificity. The informativeness of the generated text correlates strongly with coherence. The teacher’s outputs, despite containing specific details, are uninformative as a whole due to topic shifts. The 8-step student’s text has very low information content due to repetition. As NFEs increase from 16 to 128, the students’ generations become increasingly specific and informative. For example, the 64-step model constructs a detailed news-style report, and the 128-step model builds a multi-faceted story, both rich in contextual details.
@@
Text obtained with $1024$ NFEs teacher model (1/3). Perplexity=$38.53$, Entropy=$5.22$. He was mid-morning early, and I started forming a motorcade for him. He stopped every NYPD car, and I radioed back to every bicycle officer that was in there, to protect him. I did not hear you said to be involved in it. That you were involved with that campaign. I never asked anybody to test it. You did need help you leaked it, not made it for yourself. He parked and it was then that I saw that president and ended up there, speaking with a journalist. The Prince was granted anonymity and noted that the FBI was aware of the familiarity with Russia and was conducting the investigation as appropriate. In some accounts, Casey Wilson, senior counsel for the Bonden Bisson, told The Post that the President is aware of the inquiry. But Mr. Bakers testimony is the first time someone has long taken something of concern into the hands of their officials probing Trumps personal relationships with foreign officials. Former FBI regional director James Kogan, a former officer who met with President Trump, was once briefed about the inquiry from the Russian District Attorney General. Victor Atkinson, an FBI chief-officer who is retired, hacked the personal PINs and cell phone lines of the two Trump officials to conceal the information, as shown inside the countrys electronic networks. Anonymous messages can appear in our apps like Twitter and Facebook and if ones political actions resists they become blocked. The Daily Signal columnist, and former Edward Snowden contractor, last summer. According to the piece he leaked, his company was set up to log phone calls for the officials so that some of Hillarys callers would end up because Trump and businesses required a gag. I was using this same setup last week during my private conversation with Trump. I called him after this piece of political reporting and "faming," he said of Barack Obama, as he told me last May. The conversation about the leak was livestreamed at 1 p.m. this evening. CAIRCLE MEDALE, RAC.Y. Authorities a 28-year-old Burnet County receiving a call at 10:30 p.m. live across the street in Greene County reported that the man he heard was telling a neighbor to kill him. Abedoy and Terry Angelo, 24, found the house. They were charged with disturbing the peace, said Greene Police Detective Nathaniel Clay. "There is a door with a phone in a glass covering, and there is a grenade at the window," Clay said. "The caller smashed out the window, indicating it was the mans father." Hes had a similar house for a week. They were just found. Police said Angelo said he thought up the name while he was delivering. "Police determined that the most recent break in the home had been set in place," Clay said. He said assistance.
@@
Text obtained with $1024$ NFEs teacher model (2/3). Perplexity=$38.53$, Entropy=$5.22$. Montana State Supreme Court will vote Monday-2 to extend the reach of the case in Indiana until the league has their final ruling in court on November 4. The order in case reads as the first rule that the league will either settle trial court or the trial. With the lower court denying it. This means that when the ramifications stand, all players could be considered at "the game" like mayow or receive other incentives, likely a winning boost after Fridays game in Chicago. Gomez is 6-foot-6 by weight when he opens up. And probably is the future top free agent, but no players can return for taking a portion of those benefits. S. Thibodeau/AP Westwood is the standard locale to that end. Under the rules, the league must travel a minimum of four minutes per day to represent Finals 2 and 4. To every benefit from all the games, and that only players who participate play each others for each game. The league, which determines the future status of the highest free agent, will head to court in Westwood by Wednesday to avoid the headaches. The league serves as trial court in the case -- presumably with the court, which then gets to decide on any procedural cases that may arise -- before June 18, although the league says its not likely. The current clause disallows the NBA from linking any agent to the league. And prohibits it from accepting unany sponsorships that may have affected players without the approval of a director. Through its appeal, the NBA believes the Westwood will decide whether to receive any. It will update the league as it makes its decision the other day. If results are not appealed, the league ordered the court in it its ruling on March 29. It considered for an appeals court ruling on Monday, but that order has not been taken up by courts. The case seemed to already be dead Friday. After all of the courts first 18 officials handed the briefcase and could not earn a reach ruling in court itself, but it argued that it was hearing their appeal. The rule, when it is written, sends the 16-appointed back to a trial. And if they appeals the decision by following Friday to sea next Tuesday, hopefully the dispute will then be determined by the trial court. If the case is settled, the success of the Indiana incident in September could determine if the trial court can use its argument against the league and the NBA. Once both sides have signed a settlement agreement. The league says that under Indiana law the league must dismiss rulings Nov. 4 by the lower court, meaning that if it claims that a trial ruling will not be valid, it will be too late to qualify.
@@
Text obtained with $1024$ NFEs teacher model (3/3). Perplexity=$38.53$, Entropy=$5.22$. And even less because to me the LGBT community are homophobic. Were really unforgressive because generally part of us is about equality. And I guess people really do not know, if were LGBT and were not going to pay for it, wouldnt we just do it? Well, yeah, yeah, yeah, that would be great. No, I mean realized we would pay for that because then youd just have a really bad time. If something is seriously important to you and wanting to be able to participate participate in it, youd be great at that. What do you think would be of stepping up to support your local team? Is there any way your parents would have to be able to if you played to the next level. Do they feel that responsibility as parents? Yes, but I do think that when you grow up in an LGBT environment, a lot of stuff is going to happen. Thats why it has been about anything that is civil. Look, I think that its working, but obviously a family can never get through that hierarchy here because younger people are very well theyre raised, and theyre going through a lot of bad experiences around nobody that are parents. Growing up from a parent figure is just a less enjoyable environment to be able to have. We are telling the story here, so you have an answer for it. How long have you and your band have toured a college district? When were you in college? Hunsheeling Foo since I was in college. Ive never really thought about recruiting for a high school. But it wasnt just just me going onto seeing whatever Neil did or why Google just made this about it all, or it was me going into every second week. Or going into at least some draft, though, maybe thats different than knowing what this thing is. Thats why I scholarship, that made the sense. I only broke records for two years at Kansas State and Kansas, I graduated and I had many great battles afterward, so theyre always in discussion. And I have a friend recruited as a part of the Michigan teams and Michigan. David Denselmoss, former Buffalo Bills backup QB for the Eagles that coach Bruce Allen used in recruiting from Michigan for Colorado. (He was originally planned to have been a Michigan transfer.) We happened to start your press conference because you mentioned high-profile recruits but not selected in that one, either. Dolting Peps: No question. Its certainly not an uncommon one to make that mistake. But if you do ask me theres nothing I can do wrong with being among the 60 worst coaches in football.